a culture started with 5000 bacteria. after 4 hours it grew to 5500.predict how many bacteria will be after 13 hours.
Answer by stanbon(75887)

----
A culture started with 5000 bacteria. After 4 hours it grew to 5500.
A(t) = Ao*b^t
5500 = 5000*b^4
1.1 = b^4
b = 1.1^(1/4)
----
Equation:
A(t) = 5000*1.1^(t/4)
------
Predict how many bacteria will be after 13 hours.
A(13) =5000*1.1^(13/4)
-------------------
A(13) = 6815.48
Round up to 6816 bacteria
======================================
Cheers,
Stan H.